Using the following chemical equation how many grams of water are produced by reacting 59.33 g of calcium hydroxide?

2HCl + Ca(OH)2 -> 2H2O + CaCl2

1
Expert's answer
2021-02-15T02:47:37-0500

Solution:

The balanced chemical equation:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

According to the equation above: n(Ca(OH)2) = n(H2O)/2


Moles of Ca(OH)2 = Mass of Ca(OH)2 / Molar mass of Ca(OH)2

The molar mass of Ca(OH)2 is 74.093 g mol-1.

Hence,

n(Ca(OH)2) = 59.33 g / 74.093 g mol-1 = 0.80075 mol


n(H2O) = 2 × n(Ca(OH)2) = 2 × 0.80075 mol = 1.6015 mol


Moles of H2O = Mass of H2O / Molar mass of H2O

Mass of H2O = Moles of H2O × Molar mass of H2O

The molar mass of H2O is 18.0153 g mol-1.

Hence,

m(H2O) = 1.6015 mol × 18.0153 g mol-1 = 28.8515 g = 28.85 g


Answer: 28.85 grams of water (H2O) are produced.